OpenStudy (anonymous):

Medal A 1 L I.V. bag is set to infuse over 6 hours. If the drop factor is 25 gtt/mL, determine the rate of flow (in gtt/min).

OpenStudy (danjs):

ok, can you help me , , does infuse mean empty out?

OpenStudy (danjs):

i guess the drop factor is number of drops per mL... that gtt thing

OpenStudy (danjs):

I am guessing you want to figure the flow rate in number of drops/mL required to empty 1 L in 6 hrs

OpenStudy (danjs):

So the flow rate is 1 Liter per 6 Hours, to start with

OpenStudy (danjs):

\[\frac{ 1 ~~L }{6~~hrs }\]

OpenStudy (danjs):

need to use the given info, to convert the units to what you need gtt/min

OpenStudy (danjs):

\[\frac{ 1 ~~L }{6~~hrs } * \frac{ 1~hrs }{ 60~\min }*\frac{ 1000~ mL }{ 1~ L } * \frac{ 25~gtt }{ 1~mL } = \]

OpenStudy (danjs):

just arranged it so units cancel correctly, left with gtt/min

OpenStudy (anonymous):

25/1000

OpenStudy (danjs):

625/9 GTT/MIN

OpenStudy (danjs):

(1000 * 25) / (6*60)

OpenStudy (anonymous):

so the answer is 625/9?